Steering Manufacturing produces snow shovels. The selling price per snow shovel is $29.00. There is no beginning inventory. Costs involved in production are: Direct material $5.00 Direct labor 4.00 Variable manufacturing overhead 4.00 Total variable manufacturing costs per unit $13.00 Fixed manufacturing overhead per year $181,170 In addition, the company has fixed selling and administrative costs of $172,400 per year. During the year, Steering produces 54,900 snow shovels and sells 49,560 snow shovels.

How much fixed manufacturing overhead is in ending inventory under full costing?

Fixed manufacturing overhead in ending inventory $____________
Answer #1
$17,622
Statement showing Computations
Particulars Amount
Fixed manufacturing overhead per year           181,170.00
No of units produced            54,900.00
Cost per unit = 181,170/54,900                      3.30
Ending inventory in units = 54,900 - 49,560              5,340.00
Fixed manufacturing overhead in ending inventory = 5340*3.30            17,622.00